29th EACM

→ Europe/Zurich
530/R-030 - BE-EA Conference Room (CERN)

530/R-030 - BE-EA Conference Room

CERN

35
Show room on map
Silvia Schuh (CERN), Benoit Salvant (CERN), Bastien Rae (CERN)
Zoom Meeting ID
65971797648
Host
Bastien Rae
Useful links
Join via phone
Zoom URL
    • 14:00 → 14:05
      News 5m
      • Approval of the 28' EACM Minutes
      Speakers: Benoit Salvant (CERN), Silvia Schuh (CERN)
    • 14:05 → 14:10
      Action items 5m
      Speakers: Benoit Salvant (CERN), Silvia Schuh (CERN)
    • 14:10 → 14:25
      Key information from different meetings 15m
      Speakers: Bastien Rae (CERN), Benoit Salvant (CERN), Giulia Romagnoli (CERN), Natalya Kahn, Silvia Schuh (CERN)
    • 14:25 → 14:40
      Safety 15m
      Speakers: Iga Maria Krautsztrung, Isabelle Mardirossian (CERN), James Devine (CERN)
    • 14:40 → 15:00
      Coordination 20m
    • 15:00 → 15:05
      Further Lessons Learned and Opportunities for Improvement 5m
      • what must we absolutely keep doing because it works well

      • what must we stop doing

      Speaker: Michael Lazzaroni (CERN)
    • 15:05 → 15:15
      News from Experiments running during LS3 10m
      • MADMAX
        • Meeting this morning with CERN https://indico.cern.ch/event/1727421/
      • RADES
      • CMS GEM
      • Neutrino Platform
      • GIF++ : 
        • Piping work in progress
        • added 2 ODH detectors.
      • AMBER :
        • Nothing special to be reported except of an issue with our winter chiller system which is followed up by Fred from EN/CV with an maintenance of the system.
      • CLOUD
        • On-site installation work ongoing in view of CLOUD19 run start on Monday 19 October.
          • Reception of instruments. These occupy space in the entrance area next to CLOUD.
          • Installation of the instruments to the CLOUD zone. 
        • Safety:
          • CLOUD Facility ISD (Initial Safety Declaration) 2026, EDMS: 3479628
          • CLOUD19 Instruments/gases/chemicals ISD and Safety Panel, EDMS: 3479627
          • Safety inspections scheduled for 12 and 16 October.
      • ELENA/GBAR
      • ISOLDE/PUMA
      • NA61/SHINE (self-added)
        • The air in barracks got very warm after turning off chilled water circuits (28-30 deg C). Ventilation units still working. Asked EN-CV for intervention. Frederic Juban turned off the ventilation units. Temperature dropped to 21-25 deg C except for the room 887/1-A52 where it is still at 27-28. Not a big problem for us now.
        • The same thing happened to our MTPC hut (887/R-F60 + 887/1-D58). Here it is more critical as we have there fragile electronics. Together with EN-CV we turned on our old chiller located just next to 887/R-F60. Temperature being maintained now at 21-22 deg C.
        • We turned back on monitoring of atmospheric conditions in our rooms in EHN1.
    • 15:15 → 15:20
      A.O.B: NA-TARG access rights 5m